### v2

# Leveraging Multi-Modal Information to Enhance Dataset Distillation

Abstract
Dataset distillation aims to create a compact and highly representative synthetic dataset that preserves the knowledge of a larger real dataset. While existing methods primarily focus on optimizing visual representations, incorporating additional modalities and refining object-level information can significantly improve the quality of distilled datasets. In this work, we introduce two key enhancements to dataset distillation: caption-guided supervision and object-centric masking. To integrate textual information, we propose two strategies for leveraging caption features: the feature concatenation, where caption embeddings are fused with visual features at the classification stage, and caption matching, which introduces a caption-based alignment loss during training to ensure semantic coherence between real and synthetic data. Additionally, we apply segmentation masks to isolate target objects and remove background distractions, introducing two loss functions designed for object-centric learning: masked feature alignment loss and masked gradient matching loss. Comprehensive evaluations demonstrate that integrating caption-based guidance and object-centric masking enhances dataset distillation, leading to synthetic datasets that achieve superior performance on downstream tasks.
1 Introduction
Computer vision has rapidly evolved with the advent of deep learning, enabling breakthroughs in tasks such as image classification, segmentation, and object detection. These advances have been largely driven by the availability of large-scale datasets, such as ImageNet-1K, which provide the rich visual diversity necessary for training high-performance models. However, the growing reliance on massive datasets has led to significant computational, storage, and energy demands, posing challenges for efficient training, model deployment, and scalability, especially in resource-constrained environments.
Researchers have sought ways to reduce dataset sizes while preserving their training efficacy to lower storage and computational requirements. Traditional techniques, such as data pruning, coreset selection, and compression, focus on selecting or compressing informative subsets of real images. However, these methods still rely on storing real data, which can be inefficient. Dataset distillation takes a different approach, which aims to synthesize a small set of optimized samples that encapsulate the essential information of the full dataset. By distilling knowledge into a condensed form, models are expected to achieve competitive performance with significantly fewer data samples. The challenge lies in effectively capturing the complexity of high-dimensional datasets, such as ImageNet-1K, while ensuring generalization, making dataset distillation a crucial research direction for efficient and scalable deep learning.
Recent advances in dataset distillation have primarily focused on two key directions: matching-based methods [30, 27, 29, 1] and generative model priors [2, 17], both of which aim to generate synthetic datasets that closely approximate real data distributions. Matching-based methods optimize the alignment between real and synthetic images, ensuring that the distilled dataset preserves essential structural and statistical properties. These approaches have shown effectiveness in image-only datasets by directly optimizing synthetic samples against real data distributions. Meanwhile, generative model priors leverage GAN-based or diffusion models to synthesize high-quality samples, capturing complex variations in the data. While both techniques have demonstrated success in image-based tasks, they largely overlook the potential benefits of multi-modal information. Recent multi-modal methods [24, 23] have introduced image-text similarity objectives in vision language tasks. To the best of our knowledge, our approach is the first to comprehensively leverage multi-modal data, such as caption descriptions and object-centric features, from the widely-used ImageNet-1K dataset to assist the distillation process.
Building on this insight, we propose a multi-modal dataset distillation framework that integrates diverse data modalities, such as caption descriptions, segmentation masks, and bounding boxes, into the distillation process. By incorporating multi-modal cues, our approach enables the model to learn from a more comprehensive representation of the dataset, rather than relying solely on image information. Caption features provide high-level semantic understanding, masks and bounding boxes help localize important regions within images, all contributing to a more informative synthetic dataset. This integration not only enhances feature learning but also improves model robustness in downstream tasks, making dataset distillation more scalable and effective beyond conventional image-based methods.
A major challenge in incorporating multi-modal data into dataset distillation is the lack of ground truth annotations in most large-scale datasets. To address this limitation, we employ state-of-the-art methods [22, 26, 13, 7, 8] to generate captions, segmentation masks, and bounding boxes for real images. Afterwards, we propose two approaches for integrating caption features into dataset distillation. The first approach, caption feature concatenation, directly concatenates caption embeddings with visual features, allowing the model to process linguistic and visual information jointly. This fusion improves feature representation, enabling synthetic datasets to capture both semantic meaning and fine-grained visual details. The second approach, caption matching, treats captions as an additional alignment constraint rather than a fused input. During training, caption features are generated for synthetic images and matched against real image captions alongside gradient matching. This ensures that the synthetic dataset aligns not only in appearance but also in semantic content, preserving textual descriptions.
For leveraging segmentation masks, we propose masked gradient matching and masked distribution matching to enhance object-centric learning. By applying masks to both real and synthetic images, we obscure background regions, forcing the model to focus on salient object areas and preventing overfitting to irrelevant background information. The first method applies gradient matching between masked real and synthetic images while also enforcing gradient alignment on full images. The second method combines distribution matching for masked real and synthetic images with gradient matching for full images. These approaches enable the synthetic dataset to capture detailed object structures while maintaining a broader image context, enhancing robustness across downstream tasks. By integrating captions and masks into dataset distillation, our approach enriches feature learning, enhances generalization, and improves performance in downstream tasks.
Our contribution are as follows:
  1. 1.
    We integrate multi-modal information into dataset distillation by leveraging captions and segmentation masks to enhance feature representation. Due to the absence of ground-truth multi-modal annotations, we generate these annotations using pre-trained models, ensuring consistency and scalability in the distillation process.
  2. 2.
    We propose two approaches for incorporating caption features into dataset distillation: (i) Caption Concatenation: Caption features are concatenated with visual features before the classification stage, enriching semantic representations. (ii) Caption Matching: Caption features for synthetic images are generated during training, and a caption matching loss is applied to align synthetic images with real samples.
  3. 3.
    We develop two methods that utilize segmentation masks to enhance dataset distillation. We introduce mask-based gradient matching and distribution matching to optimize the learning of object-specific features while preserving semantic consistency between real and synthetic images.
  4. 4.
    We conduct extensive experiments to validate the effectiveness and generalization ability of our proposed methods. Our results demonstrate consistent performance improvements across various dataset subsets and model architectures.
2 Related work
Dataset distillation was initially introduced together with model selection [21, 18]. Subsequently, matching methods have been explored that reduce the distance between the training process on synthetic images and the same teacher model trained on the real data, such as Dataset Condensation with Gradient Matching (DC) [30, 27, 25], Distribution Matching (DM) [29, 31], Matching Training Trajectories (MTT) [1], Sequential Subset Matching [5], and feature alignment of convolutional networks [20, 14]. To improve performance, other approaches, such as factorization [10], accumulated trajectory errors [4], calibration techniques [32], and Frequency Domain utilization [16] are also proposed. Because of the high-frequency noise in pixel space, [28, 2] focus on synthesizing images in the latent space using pre-trained GAN-type generative models. Other methods address different phases of dataset distillation, such as the concept of the distillation space [9], the clustering process that selects real images for the following matching [11], and new matching metrics with mutual information [15]. Sun et al. [19] proposed cropping patches from real images and concatenating high-scoring patches to generate synthetic images. Recently, several approaches have been proposed that involve training a diffusion model as part of the dataset distillation process. Su et al. [17] trained a latent diffusion model on prototypes rather than using traditional matching-based optimization, enabling the generation of more diverse and high-quality synthetic datasets. Gu et al. [6] fine-tuned a diffusion model to generate realistic images while introducing a minimax criterion to simultaneously improve representativeness and diversity, ensuring a more balanced and informative synthetic dataset. For large-scale datasets like the full ImageNet-1K, While prior works have explored vision language methods [24, 23] on other datasets for dataset distillation. Our approach is the first to comprehensively leverage multi-modal data of the widely-used ImageNet-1K specifically for the dataset distillation process, demonstrating notable improvements over existing methods.
Figure 1: Annotations of a sample from flamingo class.
Figure 2: Overview of the Caption Combination Framework. (a) Caption Concatenation: The caption feature is integrated with the image feature before being passed through the linear layer for probability prediction, enriching the visual representation with semantic context. (b) Caption Matching: In each iteration, caption features are extracted from synthetic images and aligned with those from real images, enforcing consistency between the generated and original data representations.
3 Method
Problem introduction.
Dataset distillation refers to the process of compressing the rich information from a large real dataset into a significantly smaller synthetic dataset. The goal is to ensure that the distilled dataset retains essential characteristics, allowing models trained on it to achieve performance comparable to those trained on the full dataset in downstream tasks such as classification.
Formally, given a real dataset $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{3\times H\times W}$ denotes an image, $y_{i}\in\{0,1,2,...,C\}$ represents its corresponding class label, in a dataset with $C$ total classes and $N$ total samples. Our task is to synthesize $IPC$ images per class and form a small synthetic dataset $\mathcal{S}=\{(s_{i},y_{i})\}_{i=1}^{N_{s}}$, where $N_{s}=ipc\times C$ and $N_{s}\ll N$. The downstream tasks trained on the distilled dataset $S$ are expected to achieve strong performance, demonstrating the effectiveness of the distillation process in preserving essential information while reducing data complexity.
3.1 Multi-modal annotations generation
One of the main challenges in integrating multi-modal data is the absence of ground truth annotations in large-scale datasets, requiring automated generation methods. To address this, we utilize state-of-the-art pretrained models to generate captions, bounding boxes, and segmentation masks, following the approach introduced in InstanceDiffusion [22], a recent text-based image generation framework. Figure 1 provides an example of the generated annotations.
As a preliminary step, we generate all labels for each image to establish a broad semantic understanding. To achieve this, we employ the Recognize Anything Model (RAM) [26], a strong image-tagging model capable of detecting and assigning multiple labels to an image. RAM is trained on diverse datasets and optimized for zero-shot recognition, making it well-suited for large-scale image labeling without requiring manual annotations. By leveraging RAM, we obtain a rich set of class labels that accurately describe the objects and scenes within each image.
To generate bounding boxes for individual objects within an image, we apply Grounded Segment Anything (Grounded-SAM) [13], a model that integrates object grounding with segmentation. Grounded-SAM is particularly useful for detecting and localizing multiple objects within an image, even in cases where predefined class labels are unavailable. Once the bounding boxes are obtained, we refine the segmentation by utilizing the Segment Anything Model (SAM) [7] to generate pixel-wise masks for each detected object.
To generate descriptive captions, we employ BLIP-V2 [8], a state-of-the-art Vision-Language Model (VLM) designed for multi-modal understanding. BLIP-V2 is trained on large-scale vision-language datasets and is capable of generating coherent and contextually relevant textual descriptions for visual inputs. By feeding the cropped object instances into BLIP-V2, we obtain instance-level captions that describe each detected object. These captions are then processed using CLIP [12] to generate corresponding caption features.
3.2 Caption combination
Caption information introduces additional semantic context that enhances the quality of the distilled dataset. Captions offer a higher-level understanding of object attributes, relationships, and contextual relevance. Integrating such information can lead to more meaningful and generalizable representations in the distilled dataset. A fundamental challenge, however, is determining how to effectively integrate caption features into the distillation process without disrupting the optimization dynamics. To address this challenge, we propose two distinct strategies for incorporating caption features into dataset distillation. The first method introduces caption features at the classification stage by concatenating them with visual features before making final predictions. The second method generates caption features for synthetic images during the training distillation process and computes a caption matching loss between real and synthetic caption features to reinforce semantic consistency. Additionally, gradient matching is applied to further align the synthetic dataset with real data representations.
Caption feature concatenation.
One intuitive way to introduce caption features is by incorporating them into the classification pipeline. In this approach, we extract caption embeddings and concatenate them with the feature representations obtained from the penultimate layer of the convolutional network. This fused feature vector is then passed to the classifier, which predicts class probabilities based on both visual and textual information. The overall process is illustrated in Figure 2 a) Caption Concatenation. This method allows the model to integrate linguistic and visual representations before the final classification stage, potentially improving the discriminability of similar classes. By providing additional contextual information, caption features may help resolve ambiguities that arise when visually similar classes share overlapping features but have distinct semantic meanings.
To ensure that our method aligns with existing dataset distillation strategies, we adopt the framework of GLAD [2]. In GLAD, the distilled dataset is optimized by matching gradients between the synthetic and real datasets. Our modification extends this framework by introducing caption features at the classification stage while keeping the core distillation process unchanged. This ensures that the distilled dataset remains effective for training downstream models without requiring major modifications to the optimization pipeline. A key advantage of this approach is its computational efficiency. Since captions are generated in advance and introduced only at the classification stage, the distillation process itself remains largely unaffected. This makes the method easy to integrate into existing frameworks while still allowing models to benefit from additional textual context.
Caption matching
While the feature concatenation approach introduces textual information at the classification stage, it does not influence the underlying distilled dataset during training. To address this limitation, we propose an alternative method in which caption features are generated for synthetic images at each iteration and incorporated into the distillation process.
In this approach, as illustrated in Figure 2 b) Caption Matching, we apply pretrained BLIP-V2 [8] to generate caption features for synthetic images, and a caption matching loss is introduced as an additional optimization objective. This loss ensures that the generated images maintain semantic consistency with real images by aligning real and synthetic representations throughout the distillation process. The caption matching loss is jointly optimized with the standard gradient matching loss used in dataset distillation. Specifically, the optimization objective is formulated as follows:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{grad}+\lambda_{2}\mathcal{L}_{caption}$ (1)
where $\mathcal{L}_{grad}$ represents the traditional gradient matching loss, $\mathcal{L}_{caption}$ enforces similarity between real and synthetic captions, and $\lambda_{1},\lambda_{2}$ are weighting factors that balance the two losses. By integrating textual supervision directly into the distillation process, this approach ensures that the distilled dataset preserves both visual and linguistic characteristics. Unlike the feature concatenation method, where captions are introduced only at the classification stage, this method influences the dataset itself, potentially leading to more semantically meaningful synthetic samples.
Figure 3: Overview of the Mask Matching Framework. (a) Masked Distribution Matching: Distribution matching is applied to align the feature representations of real and synthetic images. (b) Masked Gradient Matching: gradient matching is applied, ensuring that the synthetic dataset preserves the underlying learning dynamics of the real dataset.
3.3 Object-centric alignment with masks
In conventional dataset distillation, both real and synthetic images contain a mixture of foreground objects and background elements. However, backgrounds often introduce noise and redundancy, making it difficult for distilled datasets to focus on the most informative regions. Since dataset distillation aims to generate a compact yet highly representative dataset, leveraging object-centric information can lead to better feature alignment and improved model generalization. To address this challenge, we propose utilizing segmentation masks to extract object regions from both real and synthetic images. By isolating target objects from the background, we refine the distillation process and enforce stronger alignment between synthetic and real representations. Our approach consists of two different strategies: gradient matching loss and feature-based mean squared error (MSE) loss, both applied to background-masked images.
Masked gradient matching.
We extend the idea of gradient matching to background-masked images. Traditional dataset distillation techniques optimize synthetic images by matching their gradient updates with those computed from real images. However, these methods do not explicitly account for object-centric differences, leading to potential mismatches when real and synthetic images contain varying background elements. To mitigate this issue, we compute masked gradient matching loss using only the foreground objects. Given a model parameterized by $\theta$, let $\nabla_{\theta}\mathcal{L}(x,y)$ represent the gradient of the loss function with respect to the input image $x$ and its label $y$. Our modified gradient matching objective is defined as:
$\mathcal{L}_{grad}=||\nabla_{\theta}\mathcal{L}(\hat{x}_{real},y)-\nabla_{\theta}\mathcal{L}(\hat{x}_{syn},y)||^{2}$ (2)
where both the $\hat{x}_{real}$ and $\hat{x}_{syn}$ are masked images containing only the target objects. This loss ensures that the optimization process focuses on aligning gradients derived from object features rather than being influenced by background variations. By incorporating gradient matching into the distillation process with background-masked images, we reinforce the alignment between synthetic and real data representations at a deeper level. This approach helps the model learn from the most relevant aspects of the data, leading to improved downstream performance.
Masked distribution matching.
To enforce similarity between real and synthetic representations, we compute the Mean Squared Error (MSE) loss between their extracted features. Specifically, we pass the masked real and synthetic images through the distillation network. Let $f_{\theta}(x)$ represent the feature vector obtained from an image $x$, where $\theta$ denotes the network parameters. The MSE loss is computed as follows:
$\mathcal{L}_{MSE}=\frac{1}{N_{B}}\sum_{i=1}^{N_{B}}\sum_{j=1}^{\text{IPC}}||f_{\theta}(\hat{x}_{real,i})-f_{\theta}(\hat{x}_{syn,j})||^{2}$ (3)
where $\hat{x}_{real,i}$ and $\hat{x}_{syn,j}$ represent the real and synthetic images after background masking, respectively, and $N_{B}$ is the number of images in the batch. By applying MSE loss at the feature level, we ensure that the synthetic images capture the essential object-specific representations present in the real dataset. This encourages the distilled dataset to preserve the most informative aspects of the original data while discarding irrelevant background information.
ImNet-A ImNet-B ImNet-C ImNet-D ImNet-E ImNette ImWoof ImNet-Birds ImNet-Fruits ImNet-Cats
MTT [1] 51.7±0.2 53.3±1.0 48.0±0.7 43.0±0.6 39.5±0.9 41.8±0.6 22.6±0.6 37.3±0.8 22.4±1.1 26.6±0.4
GLAD(MTT) [2] 50.7±0.4 51.9±1.3 44.9±0.4 39.9±1.7 37.6±0.7 38.7±1.6 23.4±1.1 35.8±1.4 23.1±0.4 26.0±1.1
DM [29] 39.4±1.8 40.9±1.7 39.0±1.3 30.8±0.9 27.0±0.8 30.4±2.7 20.7±1.0 26.6±2.6 20.4±1.9 20.1±1.2
GLAD(DM) [2] 41.0±1.5 42.9±1.9 39.4±0.7 33.2±1.4 30.3±1.3 32.2±1.7 21.2±1.5 27.6±1.9 21.8±1.8 22.3±1.6
DC [30] 43.2±0.6 47.2±0.7 41.3±0.7 34.3±1.5 34.9±1.5 34.2±1.7 22.5±1.0 32.0±1.5 21.0±0.9 22.0±0.6
GLAD (DC) [2] 44.1±2.4 49.2±1.1 42.0±0.6 35.6±0.9 35.8±0.9 35.4±1.2 22.3±1.1 33.8±0.9 20.7±1.1 22.6±0.8
Cap Cat (DC) 46.5±1.1 49.0±0.8 44.3±1.0 36.9±1.2 36.0±0.9 36.5±1.9 23.0±0.9 34.2±1.6 22.6±1.3 23.5±1.4
Cap Match (DC) 46.4±0.8 48.7±0.4 42.8±1.0 35.0±1.7 34.5±1.1 36.1±1.2 23.4±0.7 33.9±1.2 21.4±1.5 22.7±1.0
Masked DM (DC) 45.9±2.0 50.0±1.7 43.7±1.7 35.7±1.4 35.2±0.9 35.6±0.7 22.6±1.1 34.1±1.3 22.0±1.1 23.5±0.9
Masked DC (DC) 46.5±1.4 48.6±1.6 43.2±0.2 35.1±2.0 35.0±1.2 36.2±1.6 23.5±1.0 33.4±1.8 21.8±1.7 22.3±0.7
Table 1: Results for IPC$=1$ on subsets of ImageNet-1K at a resolution $128\times 128$. Both distillation and classification are trained using the ConvNet. DC denotes Dataset Condensation, DM denotes distribution matching, MTT denotes Matching Training Trajectories.
ImNet-A ImNet-B ImNet-C ImNet-D ImNet-E ImNette ImWoof ImNet-Birds ImNet-Fruits ImNet-Cats
DM [29] 27.2±1.2 24.4±1.1 23.0±1.4 18.4±1.7 17.7±0.9 20.6±0.7 14.5±0.9 17.8±0.8 14.5±1.1 14.0±1.1
GLAD(DM) [2] 31.6±1.4 31.3±3.9 26.9±1.2 21.5±1.0 20.4±0.8 21.9±1.1 15.2±0.9 18.2±1.0 20.4±1.6 16.1±0.7
DC [30] 38.7±4.2 38.7±1.0 33.3±1.9 26.4±1.1 27.4±0.9 28.2±1.4 17.4±1.2 28.5±1.4 20.4±1.5 19.8±0.9
GLAD(DC) [2] 41.8±1.7 42.1±1.2 35.8±1.4 28.0±0.8 29.3±1.3 31.0±1.6 17.8±1.1 29.1±1.0 22.3±1.6 21.2±1.4
Cap Cat (DC) 43.4±1.3 43.0±1.4 37.0±1.1 29.4±1.3 30.3±1.4 32.8±1.8 19.3±1.0 30.1±1.0 23.5±1.1 20.8±1.1
Cap Match (DC) 42.9±1.1 43.3±1.1 37.8±1.1 29.0±1.3 30.7±1.5 32.9±1.0 19.4±0.6 29.7±1.1 23.1±1.3 21.4±1.0
Masked DM (DC) 42.7±1.8 43.5±1.5 37.3±1.3 30.1±1.0 31.0±1.3 33.0±1.3 19.3±1.4 30.5±1.2 23.3±1.0 21.2±0.8
Masked DC (DC) 42.4±1.4 42.6±1.1 37.8±0.2 29.4±1.2 31.7±1.2 32.8±1.0 19.5±1.4 30.3±1.0 22.8±1.7 21.6±0.8
Table 2: Cross Architecture Results for IPC$=1$ on subsets of ImageNet-1K at a resolution $128\times 128$.
4 Experiments
Datasets. We evaluate our distillation network on ImageNet-1K [3]. Specifically, we test our approach on $10$ subsets of images at a resolution of $128\times 128$, with each subset containing $10$ classes. Additionally, we assess performance on $5$ subsets at a higher resolution of $256\times 256$ to further analyze the effectiveness of our method across different image scales.
Metrics. To be able to compare to the state-of-the-art, we report the classification test accuracy in all experiments.
Implementation Details. The backbone model for all experiments is GLAD [2]. The distillation process trains a ConvNet for $3000$ epochs. To comprehensively evaluate the generalization capability of the distilled dataset, we assess its performance across a diverse model pool comprising $5$ classifiers: ConvNet, ResNet18, VGG11, ViT, and AlexNet. This evaluation strategy allows us to analyze the effectiveness of the distilled data across different network architectures, ranging from convolution-based models to transformer-based vision models. To ensure statistical robustness, each distillation and classification experiment is repeated $5$ times, and we report both the mean and standard deviation of the results.
(a) Caption concatenation
(b) Caption matching
(c) Caption concatenation
(d) Masked distribution matching
(e) Masked gradient matching
Figure 4: Qualitative results of different methods. (a) Macaw from ImNet-Birds, generated using caption concatenation. (b) The same Macaw class, generated using caption matching. (c), (d), and (e) show Parachute from ImNette, where mask-based methods effectively reduce background elements.
4.1 Comparison with state-of-the-art
We evaluate our proposed method on $10$ subsets of the ImageNet-1K dataset and compare our results from both caption-based and mask-based methods to the state-of-the-art in Table 1.
Comparison of quantitative results.
To thoroughly evaluate the effectiveness of our proposed techniques, we assess two methods that integrate caption information and two methods that utilize mask-based feature learning in Table 1. The caption concatenation (Cap Cat (DC)) approach integrates caption features by concatenating them with visual features before the classification stage. This simple yet effective method improves performance by up to $9\%$ ($22.6$ vs. $20.7$ on ImNet-Fruits) across all subsets, demonstrating the benefits of leveraging semantic information from text descriptions. The caption matching (Cap Match (DC)) aligns captions features between real and synthetic images and achieves up to $5\%$ improvement across all subsets, indicating the effectiveness of multi-modal feature alignment. The masked distribution matching (Masked DM (DC)) masks out specific regions of both real and synthetic images, enforcing dataset distillation on object-centric features while reducing background noise. The results show an increase in performance of up to $6\%$ ($22.0$ vs. $20.7$ on ImNet-Fruits) across all subsets, highlighting the effectiveness of masking in improving feature generalization. The masked gradient matching (Masked DC (DC)) enhances dataset distillation by applying gradient matching between masked real and synthetic images. The method yields a performance gain of up to $5\%$ across all subsets, demonstrating that targeted gradient updates improve model learning without unnecessary computation.
Our experimental results indicate that leveraging additional multi-modal information significantly enhances dataset distillation performance. Both caption-based methods and mask-based methods contribute to improvements, but in different ways: Caption-based methods improve semantic understanding by aligning high-level textual descriptions with visual representations, ensuring that the distilled dataset preserves rich species-level features. Mask-based methods enhance object discrimination by filtering out background noise and focusing on the most informative regions. These results validate the effectiveness of incorporating multi-modal data into dataset distillation, setting a new standard for improving scalability, efficiency, and performance in large-scale dataset distillation.
Comparison of qualitative results.
Figure 4 shows the qualitative results of two caption-based and two mask-based methods at a resolution of $128\times 128$. Figure 4 (a), generated using the caption concatenation method, effectively preserves class-specific characteristics, such as distinct shapes, color patterns, feather arrangements, and complex textures. The spatial positioning of key anatomical features, such as heads, wings, and tails, appears more natural. To highlight this, we zoom in on the head of the Macaw, showcasing finer details. Figure 4 (b), produced by the caption matching method, demonstrates structural coherence as well, particularly evident in areas such as the beak of the Macaw. Figure 4 (c) serves as a baseline for comparison with mask-based methods, using caption concatenation. Figure 4 (d), generated using masked distribution matching, enhances object boundaries while suppressing background artifacts, especially in the bottom regions highlighted by the red box. The stronger separation between the foreground and background indicates that masking helps isolate and refine object representations. Figure 4 (e), generated with masked gradient matching, effectively preserves foreground objects while minimizing background distractions as well. The enhanced contrast between objects and their backgrounds suggests that the masking mechanism effectively suppresses irrelevant gradients, directing the model to focus on more discriminative object features.
(a) Caption matching
(b) Masked gradient matching
Figure 5: Qualitative results at different resolutions. (a) Rapeseed from ImNet-A, with the left image at $128\times 128$ resolution and the right image at $256\times 256$. (b) Ruddy Turnstone from ImNet-B at thet same setting.
4.2 Cross architecture results
To assess the generalization ability of our dataset distillation approach, we conduct cross-architecture evaluations. The distillation phase is trained on a ConvNet, while the classification phase is evaluated on four different architectures: ResNet18, VGG11, ViT, and AlexNet. We report the average classification accuracy across these models to measure the transferability of our synthetic dataset across diverse network structures.
In Table 2, we observe: Caption feature concatenation (Cap Cat) achieves up to $8\%$ improvement ($19.3$ vs. $17.8$ on ImWoof), showing that enriching visual representations with semantic information improves classification performance across architectures. Caption matching (Cap Match) achieves up to $9\%$ improvement ($19.4$ vs. $17.8$ on ImWoof), highlighting the effectiveness of aligning caption semantics with real images during distillation. Masked distribution matching (Masked DM) enhances performance by $8\%$ ($30.1$ vs. $28.0$ on ImNet-D), demonstrating that masking out background regions and focusing on object-centric features improves dataset quality. Masked gradient matching (Masked DC) achieves up to $10\%$ improvement ($19.5$ vs. $17.8$ on ImWoof), reinforcing that restricting gradient updates to salient object regions improves feature learning. The effectiveness of masked methods (Masked DM and Masked DC) suggests that excluding irrelevant background features leads to a more discriminative feature space, enhancing downstream classification.
4.3 Ablation study
We extend our evaluation to higher-resolution images ($256\times 256$), with results presented in Table 3. Increasing image resolution introduces additional challenges, such as greater computational complexity, more fine-grained details to capture, and increased variance in object appearances. Our approach consistently improves performance across all subsets in cross-architecture evaluations, demonstrating its effectiveness in handling high-resolution data. The caption matching method (Cap Match) demonstrates a notable improvement of $17.9\%$ ($44.1$ vs. $37.4$) on the ImNet-A subset, indicating its effectiveness in enhancing classification performance. Similarly, the masked gradient matching method (Masked DC) achieves an improvement of $6.8\%$ ($31.3$ vs. $29.3$) on the ImNet-E subset, demonstrating its effectiveness in preserving essential object features while reducing the influence of background noise. This improvement implicates that our approach has better generalization to complex images and the methods enhance the ability of the model to extract robust and transferable features.
ImNet-A ImNet-B ImNet-C ImNet-D ImNet-E
DC [30] 38.3±4.7 32.8±4.1 27.6±3.3 25.5±1.2 23.5±2.4
GLaD [2] 37.4±5.5 41.5±1.2 35.7±4.0 27.9±1.0 29.3±1.2
Cap Match (DC) 44.1±1.2 43.1±1.4 36.9±1.4 29.4±1.5 30.0±1.5
Masked DC (DC) 43.5±1.5 42.0±1.1 37.4±1.3 29.7±0.9 31.3±1.1
Cap ConvNet 44.0±1.0 47.4±0.7 41.1±0.7 33.2±0.9 33.7±1.1
Masked ConvNet 45.2±2.1 48.1±1.1 41.5±0.9 33.0±1.7 33.0±1.5
Table 3: Results for IPC$=1$ on subsets of ImageNet-1K at a resolution $256\times 256$. Cap Match and Mask DC represent the cross-architecture evaluation results. Cap ConvNet and Mask ConvNet correspond to the results where both the distillation and classification phases are conducted using the ConvNet model.
Figure 5 shows a qualitative comparison between two resolutions: $128\times 128$ and $256\times 256$. The higher-resolution images exhibit enhanced spatial coherence and finer object details compared to their lower-resolution counterparts. This improvement is particularly evident in the refined textures, such as the increased number of clouds in the sky and the more clearly defined body parts of the Ruddy Turnstone. These enhancements contribute to greater interpretability, making the synthesized images more structurally coherent.
5 Conclusion
In this work, we integrate caption-based supervision and leverage object-centric masking matching in dataset distillation. Captions provide rich semantic context that can complement visual features, and we propose two distinct approaches, caption feature combination and caption matching for incorporating them into the distillation process. Moreover, we propose two object-centric methods, masked distribution matching and masked gradient matching, for dataset distillation by leveraging images with masked backgrounds to enhance focus on target objects. By eliminating irrelevant background details, our approach ensures that models focus on critical object features, leading to more effective and robust dataset distillation. These methods enable more compact and highly informative synthetic datasets that retain key semantic and structural properties of real data, ultimately leading to improved performance in downstream tasks.
Acknowledgments
This work was supported by the State of Bavaria, the High-Tech Agenda (HTA) Bavaria and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b180dc. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) - 440719683. Support was also received from the ERC - project MIA-NORMAL 101083647. H. Reynaud is supported by the UKRI Centre for Doctoral Training AI4Health (EP / S023283/1) and Ultromics Ldt.